AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 17:18:32 2025
Author
*
:
Subject
*
:
<p>Removed tent, octagon, Neg cables.</p> <p>Disconnected and removed Baking Pi, TC08, thermocouples.</p> <p>Tested new rails, plates, fees - it fits. Need to check on the flange next.</p> <p>Will soon turn off CARME IE514 gauge to bias test the DSSDs.</p> <p>All four detectors were bias tested using the new flexi-rigid adaptor cards. Note the LK-1 jumper was removed from the adaptor cards so that 'open line' was observed between the bias lemo and the ground. Attachment 1 shows VI curves for all detectors. The maximum leakage current for all detectors appears lower than previous bias tests eg elog 400 -> temperature was ~17C today compared to ~20C for elog 400. We also performed a bias test using the old adaptor cards to see if there was any difference between the old and new cards. No significant difference was observed. For the top left detector we initally observed a significantly large leakage current (4.2 uA at 20V, 8.7 uA at 30V) which had ~1 uA fluctuations. We left for 5 minutes to see if it stabilised and the leakage current suddenly dropped to what is shown in the plot. This leakage current then appeared stable for the rest of the test. Perhaps this effect was observed by Tom but for the top right detector in elog 421?</p> <p>We plan to connect the bias as we did in the 14N+d measurement so that we can monitor and plot the leakage current over the weekend.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown